Suppose that an asteroid traveling straight toward the center of the earth were to collide with our planet at the equator and bury itself just below the surface.

1.What would have to be the mass of this asteroid, in terms of the earth's mass M , for the day to become 27.0% longer than it presently is as a result of the collision? Assume that the asteroid is very small compared to the earth and that the earth is uniform throughout.
Answer #1

T2 = T1+0.27T1 = 1.27T1

L1 = L2


I*w1 = I2W2


(2/5)*M1*R^2*2*pi/T1 = (2/5)*M2*R^2*2*pi/T2


M2 / M1 = T2 / T1


M2 = 1.27*M1


M1 + m = 1.27M1


m = 0.27 M1